OpenStudy (anonymous):

Find the exact value of tan^-1(1).

OpenStudy (anonymous):

helppp!

jimthompson5910 (jim_thompson5910):

use a calculator to enter arctan(1) to get your answer

OpenStudy (anonymous):

@jim_thompson5910 igot 90. what is it in pi form?

jimthompson5910 (jim_thompson5910):

90 degrees = 90*(pi/180) radians 90 degrees = 90pi/180 radians 90 degrees = pi/2 radians

OpenStudy (anonymous):

i got it!! thanks!!

jimthompson5910 (jim_thompson5910):

ok great, just keep that rule in mind

jimthompson5910 (jim_thompson5910):

degrees ----> radians multiply by pi/180